# Relaxed bound on performance of quantum key repeaters and secure content of generic private and independent bits

### 29 Feb 2024 v3

@@ Abstract
Quantum key repeater is the backbone of the future Quantum Internet. It is an open problem for an arbitrary mixed bipartite state shared between stations of a quantum key repeater, how much of the key can be generated between its two end-nodes. We place a novel bound on quantum key repeater rate, which uses relative entropy distance from, in general, entangled quantum states. It allows us to generalize bound on key repeaters of M. Christandl and R. Ferrara [Phys. Rev. Lett. 119, 220506]. The bound, albeit not tighter, holds for a more general class of states. In turn, we show that the repeated key of the so called key-correlated states can exceed twice the one-way distillable entanglement at most by twice the max-relative entropy of entanglement of its attacked version. We also provide a non-trivial upper bound on the amount of private randomness of a generic independent bit.
@@ I Introduction
The Quantum Internet (QI) for secure quantum communication is one of the most welcome applications of quantum information theory. In the future, Quantum Internet qubits rather than bits will be sent. However, both quantum and classical communication suffer from notorious noise in the communication channel. This is because the channel is often realized by an optical fiber with a high attenuation parameter. To overcome this, in the ”classical” Internet we got used to, the signal on the way between sender and receiver is amplified several times in the intermediate stations by copying it. This solution does not work in the case of the Quantum Internet, which transmits qubits as the signal they carry can not be amplified on the way due to the quantum no-cloning theorem Wootters and Zurek (1982); Dieks (1982). The seminal idea of quantum repeaters Briegel et al. (1998) resolves the no-cloning obstacle.
A quantum repeater is a physical realization of a protocol that allows two far apart stations to share the secure key that can be used for the one-time pad encryption. There has been tremendous effort put into building such a facility, which is now considered the backbone of the future QI Wehner et al. (2018). The most basic quantum repeater, according to the original idea of Briegel et al. (1998) consist of an intermediate station $C=C_{1}C_{2}$ between two distant stations $A$ and $B$. The stations $A$ and $C_{1}$ obtain an entangled, partially secure quantum state from some source, and so do the $B$ and $C_{2}$. The intermediate states’ role is to swap the security content so that $A$ and $B$ share a maximally entangled state. In theory, due to the seminal idea of Briegel et al. (1998) (for other architectures, see Muralidharan et al. (2016) and references therein), this can be achieved by altering two phases: (i) that of entanglement distillation, that is making the entangled states less noisy i.e. as close as possible to the singlet state
@@
In the simplest setup mentioned above, this step enhances entanglement between $A$ and $C_{1}$ station and $C_{2}$ and $B$ respectively. The second phase (ii) is entanglement swapping i.e. Żukowski et al. (1993) teleporting one subsystem $C_{1}$ of the state distilled in point (i) on $AC_{1}$ using the other singlet $C_{2}B$ to obtain a singlet between $A$ and $B$. Increasing the number of stations results, theoretically in arbitrarily distant quantum connections realized by means of singlet state.
One of the main reasons for building the Quantum Internet is the need for secure communication. Distributing pure entangled states is only one of the ways to achieve such a communication. In general, the underlying shared output state of such a secure key distribution protocol is described by the so-called private state - a state containing privacy in ideal form Horodecki et al. (2005, 2009a). A private state containing (at least) one bit of ideal key is a bipartite state on systems $ABA^{\prime}B^{\prime}$ that takes the following form when purified to the system of eavesdropper $E$:
${1\over{\sqrt{2}}}$ $\left(|00\rangle_{AB}\otimes U_{0}^{A^{\prime}B^{\prime}}\otimes\mathrm{I}_{E}|\phi\rangle_{A^{\prime}B^{\prime}E}+\right.$
$\left.|11\rangle_{AB}\otimes U_{1}^{A^{\prime}B^{\prime}}\otimes\mathrm{I}_{E}|\phi\rangle_{A^{\prime}B^{\prime}E}\right),$ (2)
@@
For this reason, the idea of quantum repeaters has been generalized in Bäuml et al. (2015) to the notion of quantum key repeater. The quantum key repeater involves the same setup of stations $AC_{1}C_{2}B$. However, the task of the protocol is to output some private state between stations $A$ and $B$ (that need not be maximally entangled in general). The target is to distribute the secure key not in the form of pure entangled states (the singlet) but in the form of general mixed quantum states that nonetheless contain ideal security.
@@
There have been established bounds on the performance of protocols distributing the key secure against quantum adversaries via the quantum key repeaters. The first obtained works for the so-called states with positive partial transposition Bäuml et al. (2015). Later it has been partially generalized to the case of the states which resemble private states in their structure Christandl and Ferrara (2017). Specifically, in Christandl and Ferrara (2017) there has been shown a significant bound on a variant of the key repeater rate $R$ for the case of two special so-called key correlated states $\rho_{AC_{1}}=\rho_{C_{2}B}\equiv\rho$. To recall this bound below we recall also the notion of key-attacked states of the key correlated states. Namely, the key correlated states, like private states, have two distinguished subsystems $\rho_{AC_{1}}\equiv\rho_{AA^{\prime}C_{1}C_{1}^{\prime}}$ where with an abuse of notation we identify system $A$ with key part system $A$ and shielding system $A^{\prime}$ (similarly for $C_{1}$). Now, when the key part systems $AC_{1}$ got measured in computational basis $\{|ij\rangle_{AC_{1}}\}$, the state becomes the key-attacked state of the key correlated state, and is denoted as $\hat{\rho}_{AA^{\prime}C_{1}C_{1}^{\prime}}$ (see Definition 1), and analogously for $\hat{\rho}_{C_{2}C_{2}^{\prime}BB^{\prime}})$. With this notions, the bound of Christandl and Ferrara (2017) applies to those key correlated states, the key attacked-state of which are separable i.e. totally insecure:
$R^{C_{1}C_{2}\rightarrow A:B}(\rho_{AC_{1}},\rho_{CB_{2}})\leq 2E_{D}^{C\rightarrow A}(\rho).$ (3)
@@
Here $E_{D}^{\rightarrow}$ is the one-way distillable entanglement Bennett et al. (1996), that is the amount of entanglement in the form of maximally entangled states that can be distilled from $\rho_{CA}$ via local quantum operations and one-way (from Charlie to Alice) classical communication operations. The $C_{1}C_{2}\rightarrow A:B$ denotes the fact that classical communication in the repeater’s protocol takes direction from $C_{1}C_{2}$ to $AB$ and later arbitrarily between $A$ and $B$ stations.
@@
In what follows, we would like to drop the assumption that the states $\rho_{AC_{1}}$ and $\rho_{C_{2}B}$ get separable after an attack on their key parts by measurement in computational basis. That is, more precisely, that their key-attacked states are separable. The reason is that checking separability for a given quantum state is a computationally hard problem (precisely NP-hard). In that sense, we broaden the class of states for which an extension of the bound of Christandl and Müller-Hermes (2017) holds. We do so by relaxing it a bit so that the obtained bound is not tighter than the previous one but naturally less stringent.
@@
In the second part of this manuscript, we consider the scenario of private randomness distillation from a bipartite state by two honest parties Yang et al. (2019). In the latter scenario, the parties share n copies of a bipartite state $\rho_{AB}$ and distil private randomness in the form of the so-called independent states. These are states from which, after local measurements, one can obtain ideal uniform bits for two parties that are private i.e. decorrelated from the system of an eavesdropper. The operations that the parties can perform in this resource theory are (i) local unitary operations and (ii) sending system via dephasing channel. In Yang et al. (2019) there were also considered cases when communication is not allowed, and maximally mixed states are accessible locally for the parties (or are disallowed). In all these four cases, the achievable rate regions were provided there, tight in most cases.
@@
Here, we study the amount of private randomness in generic local independent states. The latter states are bipartite states $\alpha_{AA^{\prime}B^{\prime}}$ possessing one bit of ideal randomness accessible via the measurement in computational basis on system $A$. The other party, holding system $B^{\prime}$ is honest in this scenario. We consider a random local independent state and prove that, given large enough systems $d_{s}$ of $A^{\prime}$ and $B^{\prime}$, in the above-mentioned scenarios, the localisable private randomness $R_{A}$ at the system of Alice is bounded as follows:
$R_{A}(\alpha_{AA^{\prime}B^{\prime}})\leq 1+\frac{1}{2\ln 2}+O(\frac{\log_{2}d_{s}}{d_{s}^{2}})$ (4)
@@
and this rate is achievable for asymptotically growing dimension of the shielding system at $A^{\prime}$, $d_{s}$, in all the four scenarios.
@@
The remainder of the manuscript is organized as follows. Section III is devoted to technical facts and definitions used throughout the rest of the manuscript. Section IV provides the bounds on key repeater rate in terms of distillable entanglement and the Renyi relative entropy of entanglement. Section V presents a simple bound on the distillable key of a random private state. In Section VI we also provide bounds on private randomness for a generic independent state. We finalize the manuscript by short discussion Section VII. The Appendix contains the bound on the distillable key, which is formulated using an arbitrary state as a proxy in the relative entropy formulas. It also contains a short description of the main facts staying behind the randomization technique used in Section V.
@@ II Main results
[…]
$R^{C_{1}C_{2}\rightarrow A:B}(\rho_{AC_{1}},\rho_{C_{2}B})\leq\inf_{\alpha\in(1,\infty)}$
$\left[{\alpha\over\alpha-1}E_{D}^{C_{1}C_{2}\rightarrow A}(\rho_{AC_{1}}\otimes\rho_{C_{2}B})+{\tilde{E}}_{\alpha}(\hat{\rho}_{AC_{1}}\otimes\hat{\rho}_{C_{2}B})\right],$ (5)
where $\hat{\rho}_{AC_{1}}$ and $\hat{\rho}_{C_{2}B}$ are the key-attacked states of the key correlated states $\rho_{AC_{1}}$ and $\rho_{C_{2}B}$ respectively (see Definition 1). Although the above bound is not tighter than the known one, it holds for a more general class of states: we do not demand the key-attacked states $\hat{\rho}$ to be separable. We note here that for the case when $\hat{\rho}$ is separable, we recover the bound (3) by taking the limit of $\alpha\rightarrow\infty$. In this case, $\tilde{E}_{\alpha}$ tends to the max-relative entropy Datta (2009a), which is zero for separable states, while the factor $\frac{\alpha}{\alpha-1}$ goes to $1$. In proving the above bound, we base on the strong converse bound on private key recently shown in Wilde et al. (2017); Das et al. (2020) (see also Christandl and Müller-Hermes (2017); Das et al. (2021) in this context). In particular, the bound given above implies, for an arbitrary key correlated state
$R^{C_{1}C_{2}\rightarrow A:B}(\rho,\rho)\leq 2E_{D}^{C_{1}\rightarrow A}(\rho)+2E_{max}(\hat{\rho}),$ (6)
@@
We further show an upper bound on the distillable key of a random private bits Horodecki et al. (2005, 2009a). These states contain ideal keys for one-time pad encryption secure against a quantum adversary who can hold their purifying system. The importance of this class of states stems from the fact that, as shown recently, they can be used in the so-called hybrid quantum key repeaters to improve the security of the Quantum Internet Sakarya et al. (2020). The first study on generic private states was done in Christandl et al. (2020). We take a different randomization procedure than the one utilized there. We base it on the fact that every private bit can be represented by, in general, not normal, operator $X$ Horodecki et al. (2009a). The latter, in turn, can be represented as $X=U\sigma$ for some unitary transformation and a state $\sigma$. We utilize techniques known from standard random matrix theory (see Appendix) to draw a random $X$ and upper bound the mutual information of the latter state, half of which upper bounds the distillable key Christandl and Winter (2004). Based on this technique via the bound of Eq. (6) we show, that a randomly chosen private bit $\gamma_{rand}$, satisfies:
$R^{C_{1}C_{2}\rightarrow A:B}(\gamma_{rand},\gamma_{rand})\leq 2E_{D}^{\rightarrow}(\gamma_{rand})+2.$ (7)
Let us recall here that a private bit has two distinguished systems. That of the key part, from which the von-Neumann measurement can directly obtain the key in the computational basis, and the system of shield, the role of which is to protect the key part from the environment. In this language, the state $U_{0}\sigma U_{0}^{\dagger}$ and $U_{1}\sigma U_{1}^{\dagger}$ are the ones appearing on its shield part, given the key of value 00 (or respectively $1$) is observed on its key part, and are called the conditional states. Any private bit $\gamma_{2}$ has distillable key at least equal to $1+\frac{1}{2}K_{D}(U_{0}\sigma U_{0}^{\dagger}\otimes U_{1}\sigma U_{1}^{\dagger})$as it is shown in Horodecki et al. (2018). From what we prove here, it turns out that the amount of key that can be drawn from the conditional states is bounded by a constant irrespective of the dimension of the shield part. More precisely, we improve on the bound from Eq. (7) by showing that a random private bit $\gamma_{rand}$ satisfies

### 09 Mar 2026 v4

# Limitations on quantum key repeaters for all key correlated states

@@ Abstract
Quantum key repeater is the backbone of the future Quantum Internet. It is an open problem to determine, for an arbitrary mixed bipartite state shared between the stations of a quantum key repeater, how much key can be generated between its two end-nodes. We place a novel bound on the quantum key repeater rate, which uses the relative entropy distance from, in general, entangled quantum states. It allows us to generalize bounds on key repeaters of M. Christandl and R. Ferrara [Phys. Rev. Lett. 119, 220506]. As in the latter article, we consider a scenario used for measurement-device-independent quantum cryptography. The derived bound, although not tighter, holds for a more general class of states, thereby avoiding the NP-hard separability problem. We show that the repeated key of the broad class of key correlated states can exceed twice the one-way distillable entanglement by at most the max-relative entropy of entanglement of its attacked version. We also provide a non-trivial upper bound on the amount of private randomness of a generic independent bit - a state containing one bit of ideal private randomness.
@@ I Introduction
The Quantum Internet (QI) for secure quantum communication is one of the most welcome applications of quantum information theory. In the future, Quantum Internet qubits rather than bits will be sent. However, both quantum and classical communication are subject to notorious noise in the communication channel. This is because the channel is often implemented using an optical fiber with a high attenuation parameter. To overcome this, in the ”classical” Internet we are used to, the signal in transit between sender and receiver is amplified several times by intermediate stations through copying. This solution does not work in the case of the Quantum Internet, which transmits qubits, as the signal they carry cannot be amplified on the way due to the quantum no-cloning theorem Wootters and Zurek (1982); Dieks (1982). The seminal idea of quantum repeaters Briegel et al. (1998) resolves the no-cloning obstacle.
A quantum repeater is a physical realization of a protocol that enables two distant stations to share a secure key that can be used for one-time pad encryption. There has been tremendous effort invested in building such a facility, which is now considered the backbone of future QI Wehner et al. (2018). The most basic quantum repeater, according to the original idea of Briegel et al. (1998), consists of an intermediate station $C=C_{1}C_{2}$ between two distant stations $A$ and $B$. The stations $A$ and $C_{1}$ receive an entangled, partially secure quantum state from a source, as do the stations $B$ and $C_{2}$. The role of intermediate states is to swap the security content, thereby enabling $A$ and $B$ to share a maximally entangled state. In theory, due to the seminal idea of Briegel et al. (1998) (for other architectures, see Muralidharan et al. (2016) and references therein), this can be achieved by altering two phases: (i) that of entanglement distillation, that is, making the entangled states less noisy i.e., as close as possible to the singlet state
@@
In the simplest setup mentioned above, this step enhances entanglement between $A$ and $C_{1}$ station and $C_{2}$ and $B$ respectively. The second phase (ii) is entanglement swapping Żukowski et al. (1993), i.e., teleporting one subsystem $C_{1}$ of the state distilled in point (i) on $AC_{1}$ using the other singlet $C_{2}B$ to obtain a singlet between $A$ and $B$. Increasing the number of stations results, theoretically, in arbitrarily distant quantum connections realized by means of the singlet state.
One of the main reasons for building the Quantum Internet is the need for secure communication. Distributing pure entangled states is only one of the ways to achieve such communication. In general, the underlying shared output state of such a secure key distribution protocol is described by the so-called private state - a state containing privacy in ideal form Horodecki et al. (2005, 2009a). A private state containing (at least) one bit of ideal key is a bipartite state on systems $ABA^{\prime}B^{\prime}$ that takes the following form when purified to the system of eavesdropper $E$:
${1\over{\sqrt{2}}}$ $\left(|00\rangle_{AB}\otimes(U_{0}^{A^{\prime}B^{\prime}}\otimes\mathrm{I}_{E})|\phi\rangle_{A^{\prime}B^{\prime}E}+\right$.
$\left.|11\rangle_{AB}\otimes(U_{1}^{A^{\prime}B^{\prime}}\otimes\mathrm{I}_{E})|\phi\rangle_{A^{\prime}B^{\prime}E}\right),$ (2)
@@
For this reason, the idea of quantum repeaters has been generalized in Bäuml et al. (2015) to the notion of quantum key repeater. The quantum key repeater uses the same station setup $AC_{1}C_{2}B$. However, the task of the protocol is to output a private state between stations $A$ and $B$ (which need not be maximally entangled in general). The goal is to distribute the secure key not in the form of pure entangled states (the singlet) but in the form of general mixed quantum states that nonetheless contain ideal security.
@@
There have been established bounds on the performance of protocols for distributing the key, secure against quantum adversaries, using quantum key repeaters. The first obtained works for the so-called states with positive partial transposition Bäuml et al. (2015). Later, it was partially generalized to the case of the states that resemble private states in their structure Christandl and Ferrara (2017). Specifically, in Christandl and Ferrara (2017), a significant bound has been shown on a variant of the key repeater rate $R$ for the case of two special so-called key correlated states $\rho_{AC_{1}}=\rho_{C_{2}B}\equiv\rho$. To recall this bound below, we also recall the notion of key-attacked states of the key correlated states. Namely, the key correlated states, like private states, have two distinguished subsystems $\rho_{AC_{1}}\equiv\rho_{AA^{\prime}C_{1}C_{1}^{\prime}}$ where, with an abuse of notation, we identify system $A$ with a key part of a system $A$ and shielding system $A^{\prime}$ (similarly for $C_{1}$). Now, when the key part systems $AC_{1}$ got measured in computational basis $\{|ij\rangle_{AC_{1}}\}$, the state becomes the key-attacked state of the key correlated state, and is denoted as $\hat{\rho}_{AA^{\prime}C_{1}C_{1}^{\prime}}$ (see Definition 1), and analogously for $\hat{\rho}_{C_{2}C_{2}^{\prime}BB^{\prime}})$. With these notions, the bound of Christandl and Ferrara (2017) applies to those key correlated states, the key attacked-state of which are separable, i.e., totally insecure:
$R^{C_{1}C_{2}\rightarrow A:B}(\rho_{AC_{1}},\rho_{C_{2}B})\leq 2E_{D}^{C_{1}\rightarrow A}(\rho_{AC_{1}}).$ (3)
@@
Here $E_{D}^{\rightarrow}$ is the one-way distillable entanglement Bennett et al. (1996), that is, the amount of entanglement in the form of maximally entangled states that can be distilled from $\rho_{CA}$ via local quantum operations and one-way (from Charlie to Alice) classical communication operations. The $C_{1}C_{2}\rightarrow A:B$ denotes the fact that classical communication in the repeater’s protocol takes direction from $C_{1}C_{2}$ to $AB$ and later arbitrarily between $A$ and $B$ stations. Note that this is precisely the scenario of public communication known in measurement-device-independent cryptography Lo et al. (2012); Braunstein and Pirandola (2012), hence the presented bound applies to this relevant quantum-cryptographic scenario (see Christandl and Müller-Hermes (2017); Azuma et al. (2016); Pirandola (2019); Das et al. (2021a)) in this context.
@@
In what follows, we would like to drop the assumption that the states $\rho_{AC_{1}}$ and $\rho_{C_{2}B}$ get separable after an attack on their key parts by measurement in the computational basis. That is, more precisely, that their key-attacked states are separable. The reason is that checking separability for a given quantum state is a computationally hard problem (precisely NP-hard). In that sense, we broaden the class of states for which an extension of the bound of Christandl and Müller-Hermes (2017) holds. We do so by relaxing it slightly so that the resulting bound is not tighter than the previous one but is naturally less stringent.
@@
In the second part of this manuscript, we consider the scenario of private randomness distillation from a bipartite state by two honest parties Yang et al. (2019). In the latter scenario, the parties share n copies of a bipartite state $\rho_{AB}$ and distill private randomness in the form of the so-called independent states. These are states from which, after local measurements, one can obtain ideal, uniformly random bits for two parties that are private, i.e., decorrelated from the eavesdropper’s system. The operations that the parties can perform in this resource theory are (i) local unitary operations and (ii) sending system via a dephasing channel. In Yang et al. (2019), cases in which communication is not allowed and maximally mixed states are locally accessible to the parties (or are disallowed) were also considered. In all four cases, the achievable rate regions were provided, and they were tight in most cases.
Here, we study the amount of private randomness in generic local independent states. The latter states are bipartite states $\alpha_{AA^{\prime}B^{\prime}}$ of the form
$\alpha_{AA^{\prime}B^{\prime}}=\sum_{i,j=0}^{1}{1\over 2}|i\rangle\langle j|_{A}\otimes U_{i}\sigma_{A^{\prime}B^{\prime}}U_{j}^{\dagger}$ (4)
@@
possessing one bit of ideal randomness accessible via the measurement in the computational basis on system $A$. The other party, holding system $B^{\prime}$ is honest in this scenario. We consider a random local independent state and prove that, given large enough systems $d_{s}$ of $A^{\prime}$ and $B^{\prime}$, in the above-mentioned scenarios, the localisable private randomness $R_{A}$ at the system of Alice is bounded as follows:
$R_{A}(\alpha_{AA^{\prime}B^{\prime}})\leq 1+\frac{1}{2\ln 2}$ (5)
@@
and this rate is achievable for asymptotically growing dimension of the shielding system at $A^{\prime}$, $d_{s}$, in all four scenarios.
@@
The remainder of the manuscript is organized as follows. Section III is devoted to technical facts and definitions used throughout the rest of the manuscript. Section IV provides the bounds on key repeater rate in terms of distillable entanglement and the Rényi relative entropy of entanglement. Section V presents a simple bound on the distillable key of a random private state. In Section VI, we also provide bounds on private randomness for a generic independent state. We finalize the manuscript with a short discussion in Section VII.
@@ II Main results
In this manuscript, we develop a relaxed bound on quantum key repeater rate, which holds for the key-correlated states without the assumption that the $\hat{\rho}$ is separable. This relaxation is essential, as checking separability is, in general, an NP-hard problem Gurvits (2003). We provide a bound on the key repeater rate using the relative entropy distance to arbitrary states rather than to separable ones. The freedom in choosing a state in the relative entropy distance is, however, compensated by another relative-entropy-based term (the sandwich Rényi relative entropy of entanglement $\tilde{E}_{\alpha}$ Hayashi (2017)) and a pre-factor $\alpha$ from the interval $(1,+\infty)$. It reads the following upperboundbound on $R^{C_{1}C_{2}\rightarrow A:B}(\rho_{AC_{1}},\rho_{C_{2}B})$:
$\inf_{\alpha\in(1,\infty)}\left[{\alpha\over\alpha-1}E_{D}^{C_{1}C_{2}\rightarrow A}(\rho_{AC_{1}}\otimes\rho_{C_{2}B})\right$.
$\left.\qquad+\min\{{\tilde{E}}_{\alpha}(\hat{\rho}_{AC_{1}}),{\tilde{E}}_{\alpha}(\hat{\rho}_{C_{2}B})\}\right],$ (6)
where $\hat{\rho}_{AC_{1}}$ and $\hat{\rho}_{C_{2}B}$ are the key-attacked states of the key correlated states $\rho_{AC_{1}}$ and $\rho_{C_{2}B}$ respectively (see Definition 1). Although the above bound is not tighter than the known one, it holds for a more general class of states: we do not demand the key-attacked states $\hat{\rho}$ to be separable. We note here that for the case when $\hat{\rho}$ is separable, we recover the bound (3) by taking the limit of $\alpha\rightarrow\infty$. In this case, $\tilde{E}_{\alpha}$ tends to the max-relative entropy Datta (2009a), which is zero for separable states, while the factor $\frac{\alpha}{\alpha-1}$ goes to $1$. In proving the above bound, we base on the strong converse bound on private key recently shown in Wilde et al. (2017); Das et al. (2020) (see also Christandl and Müller-Hermes (2017); Das et al. (2021b) in this context). In particular, the bound given above implies, for an arbitrary key correlated state
$R^{C_{1}C_{2}\rightarrow A:B}(\rho_{AC_{1}},\rho_{C_{2}B})\leq 2E_{D}^{C_{1}\rightarrow A}(\rho_{AC_{1}})+E_{max}(\hat{\rho}_{AC_{1}}),$ (7)
@@
We further show an upper bound on the distillable key of a random private bits Horodecki et al. (2005, 2009a). These states contain ideal keys for one-time pad encryption, secure against a quantum adversary who can hold their purifying system. The importance of this class of states stems from the fact that, as recently shown, they can be used in the so-called hybrid quantum key repeaters to enhance the security of the Quantum Internet Sakarya et al. (2020). The first study on generic private states was done in Christandl et al. (2020). We take a different randomization procedure than the one utilized there. We base it on the fact that every private bit can be represented by, in general, not normal, operator $X$ Horodecki et al. (2009a). The latter, in turn, can be represented as $X=U\sigma$ for some unitary transformation and a state $\sigma$. We utilize techniques known from standard random matrix theory (see Appendix) to draw a random $X$ and upper bound the mutual information of the latter state, half of which upper bounds the distillable key Christandl and Winter (2004). Based on this technique via the bound of Eq. (7) we show, that a randomly chosen private bit $\gamma_{rand}$, satisfies:
$R^{C_{1}C_{2}\rightarrow A:B}(\gamma_{rand},\gamma_{rand})\leq 2E_{D}^{\rightarrow}(\gamma_{rand})+1.$ (8)
Recall that a private bit has two distinguished systems. That of the key part, from which the von-Neumann measurement can directly obtain the key in the computational basis, and the system of shield, the role of which is to protect the key part from the environment. In this language, the state $U_{0}\sigma U_{0}^{\dagger}$ and $U_{1}\sigma U_{1}^{\dagger}$ are the ones appearing on its shield part, given the key of value 0 (or respectively $1$) is observed on its key part, and are called the conditional states. Any private bit $\gamma_{2}$ has distillable key at least equal to $1+\frac{1}{2}K_{D}(U_{0}\sigma U_{0}^{\dagger}\otimes U_{1}\sigma U_{1}^{\dagger})$ as it is shown in Horodecki et al. (2018). From what we prove here, it turns out that the amount of key that can be drawn from the conditional states is bounded by a constant irrespective of the dimension of the shield part. More precisely, we derive a bound on $R^{C_{1}C_{2}\rightarrow A:B}(\gamma_{rand},\gamma_{rand})$ that is independent from Eq. (8), by showing that a random private bit $\gamma_{rand}$ satisfies